Institut for Matematik og Datalogi Odense Universitet March 11, 1996

JFB

DM11 – F96 – Lecture 6

Lecture, March 6

We covered section 3.3 and structural induction. We skipped sections 3.4 and 3.5, since they are covered in other courses. We began on chapter 4, finishing up through the inclusion-exclusion principle in section 4.1.

Lecture, March 13

We will cover sections 4.1 through 4.3 of chapter 4.

Lecture, March 20

We will cover sections 4.4 through 4.7 of chapter 4.

Problems to be discussed in week 12

12, 18a, 18e, 18f, 18g, 18h, 22, 33, 36, 38, 42, 48 from section 4.1. 38 is probably the hardest. 10, 20, 32, 34 from section 4.2. If you want to think about an interesting problem, try to come up with a more efficient algorithm than the one given in the textbook from problem 19 in section 4.2. Note: You may turn in number 34 from section 4.2 to your instructor for comments.